2 trains are moving with velocities v1=10m/s and v2=20m/s on the same track in opp dirns. After tbe application of brakes if their retarding rates are 2m/s^2 and 1m/s^2 respectively then the minimum distance of separation to avoid collision is?

Grade:

1 Answers

Vasantha Kumari
10 years ago

Train A:

Moves with velocity 10 m/s and retards at 2m/s^2,

Hence, substituting values in v^2 = u^2 –2as

0 = 100 – 2 x2x s

Or s= 25 m

Train B:

Moves with velocity 20 m/s and retards at 1m/s^2,

Hence, substituting values in v’^2 = u’^2 –2as’

0 = 400 – 2 x1x s’

Or s= 200 m

So, the minimum distance of separation to avoid collision = 200+25 =225 m

Thanks & Regards,

Vasantha Sivaraj,

askIITians faculty